OpenStudy (anonymous):

Solve the following equation: 6y – 20 = 2y – 4.

OpenStudy (lgbasallote):

subtract 2y from both sides first \[\implies 6y - 20 - 2y = 2y - 4 - 2y\] combine like terms \[\implies 4y - 20 = -4\] now add 20 to both sides \[\implies 4y - 20 + 20 = -4 + 20\] combine like terms \[\implies 4y = 16\] now divide both sides by 4 \[\implies \frac{4y}{4} = \frac{16}{4}\] simplify \[\implies y = 4\] does that help?